OpenStudy (anonymous):

solve by the substitute method 5x+5y=5 -9x+y=31

OpenStudy (anonymous):

You have the two equations 5x + 5y = 5 y - 9x = 31 You would one to use the second equation because there is a coefficient of 1 for y, which simplify the substitution method. y - 9x = 31 Move everything over from the side where y is. y = 9x + 31 Take our first equation and plug our re-arranged equation in to y 5x + 5y = 5 5x + 5(9x + 31) = 5 Multiply 5 to what is inside the parenthesis 5x + 45x + 155 = 5 Combine like terms 50x + 155 = 5 Subtract 155 on both sides 50x = -150 Divide by 50 on both sides x = -3 We now need to take one of the equations and plug -3 in for x; it doesn't matter which equation it is, so we will use the first one. 5x + 5y = 5 Plug in -3 for x 5(-3) + 5y = 5 Multiply -15 + 5y = 5 Add 15 on both sides 5y = 20 Divide by 5 on both sides y = 4 x is equal to -3 and y is equal to 4